# Combinatorics and Hodge theory of degenerations of abelian varieties: A survey of the Mumford construction

### v2

@@ 1. Introduction
In the case of a maximal degeneration, the intermediate cover $Y^{*}$ of the punctured family admits an analytic open embedding $Y^{*}\hookrightarrow({\mathbb{C}}^{*})^{g}\times{\mathbb{C}}^{*}$ into an algebraic torus, with the map to $\Delta^{*}$ given by the projection to the second factor. Thus, by the theory of toric varieties, it is possible to extend $Y^{*}\to\Delta^{*}$ to a family $Y\to\Delta$. More precisely, we take a toric extension of $({\mathbb{C}}^{*})^{g}\times{\mathbb{C}}^{*}$ for which the fiber-preserving action of ${\rm gr}_{0}^{W}V_{\mathbb{Z}}$ extends to an action on the central fiber $Y_{0}$. The quotient of $Y\to\Delta$ by the extended action of ${\rm gr}_{0}^{W}V_{\mathbb{Z}}$ produces a model $X\to\Delta$ of the degeneration.
@@ 1.1. Contents
We continue in Section 4 with a more detailed analysis of Mumford constructions associated to regular matroids (Constr. 4.16). A matroid $\underline{R}$ is a collection of subsets of a finite set $E$, encoding the notion of linear independence, of a set of vectors in a vector space (Def. 4.1). An embedding $E\hookrightarrow{\mathbb{F}}^{n}$ into a vector space over a field $\mathbb{F}$}, realizing this collection of independent sets, is a realization of $\underline{R}$. Matroids which admit a realization over any field are regular (Def. 4.2), though ${\mathbb{F}}_{2}$ and ${\mathbb{F}}_{3}$ suffice.
@@
Finally, we review the work of Alexeev, Nakamura, and Namikawa on the extension of the universal family ${\mathcal{X}}_{g}\to{\mathcal{A}}_{g}$ over the toroidal compactification $\overline{{\mathcal{A}}}_{g}^{\rm vor}$ associated to a distinguished fan ${\mathfrak{F}}_{\rm vor}$ (Defs. 6.1, 6.4), whose support is the cone ${\mathcal{P}}_{g}^{+}$ of positive semi-definite $g\times g$ matrices with rational null space.
@@ 1.2. Index of constructions
    Via fans, over the toroidal extension ${\mathcal{A}}_{g}\hookrightarrow{\mathcal{A}}_{g}^{{\mathbb{B}}}$ associated to a rational polyhedral cone ${\mathbb{B}}={\mathbb{R}}_{\geq 0}\{B_{1},\dots,B_{k}\}\subset{\mathcal{P}}_{g}^{+}$ in the space ${\mathcal{P}}_{g}^{+}$.
@@
An extensive collection of examples (Exs. 3.30, 3.33, 3.34, 3.38, 4.13, 4.14, 4.20, 4.21, 5.16, 6.7), with figures, is also provided in the text, see especially Section 3.5. The first of these (Ex. 3.30) is the prerequisite ur-example of the Mumford construction: the Tate curve, i.e. the extension of the family ${\mathbb{C}}^{*}/u^{\mathbb{Z}}\to\Delta_{u}^{*}$ of elliptic curves by an irreducible nodal curve.
1.3. Acknowledgements
PE thanks Valery Alexeev for useful discussions.
@@ 2.1. Algebraic and analytic stacks
By a DM algebraic stack, or simply DM stack, we will mean a separated Deligne–Mumford stack of finite type over $\mathbb{C}$}. Similarly, a DM analytic stack will be a separated Deligne–Mumford analytic stack $X$ in the sense of [60, Def. 5.2]. Thus, $X$ is a stack on the site of complex analytic spaces such that the diagonal is representable and finite and there exists an analytic space $Y$ and a surjective étale morphism $Y\to X$. It follows that $X$ is locally modeled as a finite quotient of an analytic space, see [60, Prop. 5.4], and that the analytification of a DM algebraic stack is a DM analytic stack, see [60, Lem. 5.5].
@@ 2.2. Principally polarized abelian varieties
A $\mathbb{Z}$}-polarized Hodge structure $(V_{\mathbb{Z}},H^{\bullet,\bullet},L)$ of weight $k$ is a $\mathbb{Z}$}-module $V_{\mathbb{Z}}$ together with an integral, non-degenerate, $(-1)^{w}$-symmetric bilinear form $L\colon V_{\mathbb{Z}}\otimes V_{\mathbb{Z}}\to{\mathbb{Z}}$, and a Hodge decomposition
@@
Consider the Lagrangian Grassmannian ${\rm LGr}(V_{\mathbb{C}},L)$. It is the projective flag variety of isotropic $g$-dimensional subspaces of $V_{\mathbb{C}}$. The polarized weight 1-1 Hodge structures on $(V_{\mathbb{Z}},L)$ with a Hodge decomposition $V_{\mathbb{C}}=H^{-1,0}\oplus H^{0,-1}$ form an analytic open subset of ${\rm LGr}(V_{\mathbb{C}},L)$, given by
@@
The condition that $H^{-1,0}$ is Lagrangian for $L$ amounts to the property that $L\in H^{1,1}(X)$ is a Hodge class. Hence $L$ determines a holomorphic line bundle ${\mathcal{L}}\to X$, unique up to translation by ${\rm Pic}^{0}(X)$. Finally, the condition $iL(\bar{v},v)>0$ ensures that $\mathcal{L}$} is ample, and so in fact, $X$ is an abelian variety (i.e. projective).
@@ 2.3. Degenerations of PPAVs
Let $f\colon(X,L)\to\Delta^{k}$ be a degeneration of PPAVs of dimension $g$ over a polydisk $\Delta^{k}$ with coordinates $u_{1},\dotsc,u_{k}$, such that the discriminant locus is the union of the coordinate hyperplanes $V(u_{i})=\{u_{i}=0\}$, for $i=1,\dots,k$. Fix a base point $t\in(\Delta^{*})^{k}$ and let $V_{\mathbb{Z}}\coloneqq H_{1}(X_{t},{\mathbb{Z}})$. Suppose that the monodromy transformation $T_{i}\colon V_{\mathbb{Z}}\to V_{\mathbb{Z}}$ associated to the simple loop $\gamma_{i}\in\pi_{1}((\Delta^{*})^{k},t)\simeq{\mathbb{Z}}^{k}$ is unipotent—for instance, by a result of Clemens [19, Thm. 7.36], this holds if the general fiber over $V(u_{i})$ has reduced normal crossings.
@@
Any collection of commuting unipotent matrices $T_{i}\in{\rm Sp}_{2g}({\mathbb{Z}})$ can be simultaneously conjugated into the standard maximal unipotent subgroup $U_{\mathbb{Z}}$. Thus, choosing a basis appropriately, we may assume that the monodromies $T_{i}$ of the degeneration $f$ are all of the form
@@
This provides a coordinate-free definition of the matrices $B_{i}\in{\rm Sym}_{g\times g}({\mathbb{Z}})$ from above.
@@
is the monodromy cone associated to the degeneration $f\colon(X,L)\to\Delta^{k}$, or more generally, family of PPAVs $f^{\ast}\colon(X^{\ast},L^{\ast})\to(\Delta^{\ast})^{k}$.
@@
Note that $\widetilde{\Psi}$ descends because the $(\tau_{1}B_{1}+\cdots+\tau_{k}B_{k})$ term cancels the equivariance of $\widetilde{\Phi}$, and so makes the map invariant under the action of ${\mathbb{Z}}^{k}$. We now recall Schmid’s multivariable nilpotent orbit theorem [56, Thm. 4.12], applied to our setting:
@@
${\mathcal{P}}_{g}^{+}=\textstyle{\mathcal{P}}_{g}\sqcup\bigsqcup_{V_{1}}{\mathcal{P}}_{g-1}\sqcup\bigsqcup_{V_{2}}{\mathcal{P}}_{g-2}\sqcup\cdots\sqcup{\mathcal{P}}_{0}$
where the $V_{i}\subset{\mathbb{Z}}^{g}$ range over all primitive integral sublattices of ${\mathbb{Z}}^{g}$ of codimension $i$, and the relevant copy of ${\mathcal{P}}_{g-i}$ is the cone of positive-definite bilinear forms on ${\mathbb{R}}^{g}/(V_{i}\otimes{\mathbb{R}})$. See Figure 22 to visualize the projectivization of ${\mathcal{P}}_{2}^{+}$, which is a cusped hyperbolic disk.
@@
We have that $W_{-2}^{\perp}=W_{-1}$ where the perpendicular is taken with respect to $L$. Thus $L$ descends to a unimodular symplectic form on $W_{-1}/W_{-2}$. In fact, by [56, Thm. 6.16], we have the following fundamental theorem on the existence of the limit MHS:
@@
with the $10$ singularities given by the $S_{6}$-orbit of the point $1:1:1:-1:-1:-1]\in{\mathbb{P}}^{5}$. Then, the universal deformation (whose existence follows from [29, Cor. 4.2], for example)
@@
see [33, Sec. 7.1.10], and generate a primitive integral Lagrangian subspace $W_{-2}\subset H_{3}(Y_{t},{\mathbb{Z}}(-1))$ of rank $5$. The Dehn twist about the three-sphere $\gamma_{i}$ gives the formula
@@ 2.4. Toric varieties
Then the gluing $Y({\mathfrak{F}})\coloneqq\bigcup_{\tau\in{\mathfrak{F}}}Y(\tau)$ of these affine schemes along the natural open immersions corresponding face inclusions gives the toric variety $Y({\mathfrak{F}})$, see [30, Sec. 1.4].
@@
As in the usual theory of toric varieties, the torus orbits of dimension $d$ in $Y({\mathfrak{F}})$, isomorphic necessarily to $({\mathbb{C}}^{*})^{d}$, correspond bijectively to cones of codimension $d$ in $\mathfrak{F}$}, see [30, Sec. 3.1].
@@
The toric variety $Y=Y_{H}$ associated to the polytope $H$ is the union of ${\operatorname{Spec}}\,{\mathbb{C}}[{\bf M}_{F}]$ ranging over all faces, see also [30, Sec. 1.5]. If, furthermore, all faces of $H$ are integral polytopes, then there is a canonically defined torus-equivariant holomorphic line bundle $\mathcal{L}$} on $Y$, given by gluing together line bundles on each affine chart ${\operatorname{Spec}}\,{\mathbb{C}}[{\bf M}_{F}]$, in a manner which locally agrees with the recipe in [30, Sec. 3].
@@ 2.5. Toroidal extensions
A fan $\mathfrak{F}$} for ${\mathcal{A}}_{g}$ is a rational polyhedral decomposition whose support is contained in ${\mathcal{P}}_{g}^{+}$, for which $\mathfrak{F}$} is $\operatorname{GL}_{g}({\mathbb{Z}})$-invariant under the action $B\mapsto ABA^{T}$, with finitely many orbits of cones.
@@
$\overline{E}\colon U_{\mathbb{Z}}\backslash{\mathcal{H}}_{g}\hookrightarrow U_{\mathbb{Z}}\backslash{\rm Sym}_{g\times g}({\mathbb{C}})\simeq\operatorname{Sym}_{g\times g}({\mathbb{C}}^{\ast})\simeq{\rm Sym}_{g\times g}({\mathbb{Z}})\otimes{\mathbb{C}}^{*}\simeq U_{\mathbb{Z}}\otimes{\mathbb{C}}^{*}.$
@@
${\mathcal{A}}_{g}\hookrightarrow{\mathcal{A}}_{g}^{{\mathfrak{F}}}$
[…]
@@
Next, it follows from the exponential convergence (2) of $\Phi$ towards $\Phi_{\rm nilp}$ (in the invariant metric on ${\mathcal{H}}_{g}$) that $\Phi\colon(\Delta^{*})^{k}\to{\mathcal{A}}_{g}$ admits a continuous extension $\Delta^{k}\to{\mathcal{A}}_{g}^{{\mathbb{B}}}$. Since $\Delta^{k}$ is normal, Riemann’s removable singularities theorem implies that this continuous extension is holomorphic. It is unique because the toroidal extension is a separated analytic space. ∎
More generally, the same construction applies to any $\operatorname{GL}_{g}({\mathbb{Z}})$-invariant fan $\mathfrak{F}$}, and the resulting toroidal extension is relatively proper over the 0-dimensional cusp of the Baily-Borel [11] compactification $\overline{{\mathcal{A}}}_{g}^{\rm BB}={\mathcal{A}}_{g}\sqcup{\mathcal{A}}_{g-1}\sqcup\cdots\sqcup{\mathcal{A}}_{1}\sqcup{\mathcal{A}}_{0}$ if and only if ${\rm Supp}({\mathfrak{F}})={\mathcal{P}}_{g}^{+}$. Then, the construction can be extended over the intermediate-dimensional strata of the Baily-Borel compactification, to produce a so-called toroidal compactification
${\mathcal{A}}_{g}\hookrightarrow\overline{{\mathcal{A}}}_{g}^{{\mathfrak{F}}}.$
@@
Then, a toroidal extension of $\widetilde{{\mathcal{A}}}_{g}$ at this Baily–Borel cusp has the “advantage” that a fan need only be $\Gamma(n)$-invariant. Furthermore, the fundamental domain for the action of $\Gamma(n)$ is larger than the fundamental domain for that action of $\operatorname{GL}_{g}({\mathbb{Z}})$. In particular, given any polyhedral cone ${\mathbb{B}}\subset{\mathcal{P}}_{g}^{+}$, there exists some $n$ (depending on $\mathbb{B}$}) for which $\mathbb{B}$} embeds into the quotient $\Gamma(n)\backslash{\mathcal{P}}_{g}^{+}$. Then the above discussion applies at this cusp, since $\Gamma(n)\cdot{\mathbb{B}}$ now defines a fan.
@@
Finally, to verify that the monodromy representation of $f^{*}$ is as specified, consider a $1$-parameter family $\Delta^{*}\to Y^{*}$ which extends to a map $\Delta\to Y$ sending $0\mapsto 0$. Then the monodromy over $\Delta^{*}$ can be computed in the étale-local model as the monodromy of the family of PPAVs over a co-character. Such co-characters correspond to lattice points, in $\tau\cap{\mathbb{Z}}^{k}$. Since $f^{*}\colon X^{*}\to Y^{*}$ is pulled back from $V^{*}$, the monodromy representation is pulled back along the morphism of cocharacter lattices, given by the inclusion $\Lambda\hookrightarrow{\rm Sym}_{g\times g}({\mathbb{Z}})$ which, in particular, restricts $\phi\colon{\mathbb{Z}}^{k}\to{\rm Sym}_{g\times g}({\mathbb{Z}})$. Finally, it suffices then to observe that the monodromy over the cocharacter $B\otimes{\mathbb{C}}^{*}\subset{\rm Sym}_{g\times g}({\mathbb{Z}})\otimes{\mathbb{C}}^{*}$ is canonically identified with $B\in{\mathcal{P}}_{g}^{+}\cap{\rm Sym}_{g\times g}({\mathbb{Z}})$, see e.g. (10).

### v3

@@ 1. Introduction
In the case of a maximal degeneration, the intermediate cover $Y^{*}$ of the punctured family admits an analytic open embedding $Y^{*}\hookrightarrow({\mathbb{C}}^{*})^{g}\times{\mathbb{C}}^{*}$ into an algebraic torus, with the map to $\Delta^{*}$ given by the projection to the second factor. Thus, by the theory of toric varieties, it is possible to extend $Y^{*}\to\Delta^{*}$ to a family $Y\to\Delta$. More precisely, we take a toric extension of $({\mathbb{C}}^{*})^{g}\times{\mathbb{C}}^{*}$ for which the fiber-preserving action of ${\rm gr}_{0}^{W}V_{\mathbb{Z}}$ extends to an action on the central fiber $Y_{0}$. The quotient of $Y\to\Delta$ by the extended action of ${\rm gr}_{0}^{W}V_{\mathbb{Z}}$ produces a new model of the degeneration with particularly nice properties, e.g. toroidal singularities.
@@ 1.1. Contents
We continue in Section 4 with a more detailed analysis of Mumford constructions associated to regular matroids (Constr. 4.16). A matroid $\underline{R}$ is a collection of subsets of a finite set $E$, encoding the notion of linear independence of a set of vectors in a vector space (Def. 4.1). An embedding $E\hookrightarrow{\mathbb{F}}^{g}$ into a vector space over a field $\mathbb{F}$}, realizing this collection of independent sets, is a realization of $\underline{R}$. Matroids which admit a realization over any field are regular (Def. 4.2), though ${\mathbb{F}}_{2}$ and ${\mathbb{F}}_{3}$ suffice.
@@
Finally, we review the work of Alexeev, Nakamura, Namikawa and Faltings–Chai on the extension of the universal family ${\mathcal{X}}_{g}\to{\mathcal{A}}_{g}$ over the toroidal compactification $\overline{{\mathcal{A}}}_{g}^{\rm vor}$ associated to a distinguished fan ${\mathfrak{F}}_{\rm vor}$ (Defs. 6.1, 6.4), whose support is the cone ${\mathcal{P}}_{g}^{+}$ of positive semi-definite $g\times g$ matrices with rational null space.
@@ 1.2. Index of constructions
    Via fans, over the toroidal extension ${\mathcal{A}}_{g}\hookrightarrow{\mathcal{A}}_{g}^{{\mathbb{B}}}$ associated to a rational polyhedral cone ${\mathbb{B}}={\mathbb{R}}_{\geq 0}\{B_{1},\dots,B_{k}\}\subset{\mathcal{P}}_{g}^{+}$.
@@
An extensive collection of examples (Exs. 3.31, 3.34, 3.35, 3.40, 4.13, 4.14, 4.20, 4.21, 5.16, 6.7), with figures, is also provided in the text, see especially Section 3.5. The first of these (Ex. 3.31) is the prerequisite ur-example of the Mumford construction: the Tate curve, i.e. the extension of the family ${\mathbb{C}}^{*}/u^{\mathbb{Z}}\to\Delta_{u}^{*}$ of elliptic curves by an irreducible nodal curve.
1.3. Acknowledgments
We thank Valery Alexeev, Younghan Bae, David Holmes, and Mirko Mauri for useful discussions and comments.
@@ 2.1. Algebraic and analytic stacks
By a DM algebraic stack, or simply DM stack, we will mean a separated Deligne–Mumford stack of finite type over $\mathbb{C}$}. Similarly, a DM analytic stack will be a separated Deligne–Mumford analytic stack $X$ in the sense of [64, Def. 5.2]. Thus, $X$ is a stack on the site of complex analytic spaces such that the diagonal is representable and finite and there exists an analytic space $Y$ and a surjective étale morphism $Y\to X$. It follows that $X$ is locally modeled as a finite quotient of an analytic space, see [64, Prop. 5.4], and that the analytification of a DM algebraic stack is a DM analytic stack, see [64, Lem. 5.5].
@@ 2.2. Principally polarized abelian varieties
A $\mathbb{Z}$}-polarized Hodge structure $(V_{\mathbb{Z}},H^{\bullet,\bullet},L)$ of weight $k$ is a $\mathbb{Z}$}-module $V_{\mathbb{Z}}$ together with an integral, non-degenerate, $(-1)^{k}$-symmetric bilinear form $L\colon V_{\mathbb{Z}}\otimes V_{\mathbb{Z}}\to{\mathbb{Z}}$, and a Hodge decomposition
@@
Let $(V_{\mathbb{Z}},L)$ be a unimodular symplectic lattice, and consider the Lagrangian Grassmannian ${\rm LGr}(V_{\mathbb{C}},L)$. It is the projective flag variety of isotropic $g$-dimensional subspaces of $V_{\mathbb{C}}$. The polarized weight 1-1 Hodge structures on $(V_{\mathbb{Z}},L)$ with a Hodge decomposition of the form $V_{\mathbb{C}}=H^{-1,0}\oplus H^{0,-1}$ define an analytic open subset of ${\rm LGr}(V_{\mathbb{C}},L)$, given by
@@
The condition that $H^{-1,0}$ is Lagrangian for $L$ amounts to the property that $L\in H^{1,1}(X)$ is a Hodge class. Hence $L$ determines a holomorphic line bundle ${\mathcal{L}}\to X$, unique up to translation by ${\rm Pic}^{0}(X)$. Finally, the condition $iL(\bar{v},v)>0$ ensures that any lift $\mathcal{L}$} is ample, and so in fact, $X$ is an abelian variety (i.e. projective).
@@ 2.3. Degenerations of PPAVs
Let $f\colon(X,L)\to\Delta^{k}$ be a degeneration of PPAVs of dimension $g$ over a polydisk $\Delta^{k}$ with coordinates $u_{1},\dotsc,u_{k}$, such that the discriminant locus is the union of the coordinate hyperplanes $V(u_{i})=\{u_{i}=0\}$, for $i=1,\dots,k$. Fix a base point $t\in(\Delta^{*})^{k}$ and let $V_{\mathbb{Z}}\coloneqq H_{1}(X_{t},{\mathbb{Z}})$. Suppose that the monodromy transformation $T_{i}\colon V_{\mathbb{Z}}\to V_{\mathbb{Z}}$ associated to the simple, oriented loop $\gamma_{i}\in\pi_{1}((\Delta^{*})^{k},t)\simeq{\mathbb{Z}}^{k}$ is unipotent—for instance, by a result of Clemens [19, Thm. 7.36], this holds if the general fiber over $V(u_{i})$ has reduced normal crossings.
@@
The above weight filtration may also be described as follows:
(2) $\textstyle W_{-2}=\left(\sum_{i=1}^{k}{\rm im}\,N_{i}\right)^{\rm sat}\quad\text{and}\quad W_{-1}=\bigcap_{i=1}^{k}\ker(N_{i}).$
Indeed, by the saturatedness of the above filtration, it suffices to prove this for rational, and, in fact, for real coefficients. The inclusions ${\rm im}(\sum N_{i})\subset\sum{\rm im}\,N_{i}$ and $\bigcap\ker(N_{i})\subset\ker(\sum N_{i})$ are clear. To prove the converse, let $(r_{1},\dots,r_{k})\in({\mathbb{R}}_{>0})^{k}$ and note that ${\rm im}(\sum r_{i}N_{i})$ and $\ker(\sum r_{i}N_{i})$ do not depend on the choice of $r_{i}>0$, see [16, Thm. 3.3]. The inclusions in question thus follow from a limit argument where $r_{j}=1$ and $r_{i}\to 0$ for $i\neq j$, applied to ${\rm im}(\sum r_{i}N_{i})\subset{\rm im}(\sum N_{i})$ and $\ker(\sum N_{i})\subset\ker(\sum r_{i}N_{i})$.
@@
The collection of commuting unipotent matrices $T_{i}\in{\rm Sp}_{2g}({\mathbb{Z}})$ can be simultaneously conjugated into the unipotent subgroup $U_{\mathbb{Z}}$ as they fix a coisotropic space (given by $W_{-1}$) and hence fix a Lagrangian subspace. Thus, choosing a basis appropriately, we may assume that the monodromies $T_{i}$ are all of the form
@@
This provides a coordinate-free definition of the matrices $B_{i}$ from above. Implicit in the above definition is the claim that $N_{i}$ contains $W_{-1}V_{\mathbb{Z}}$ in its kernel, which follows from (2).
@@
is the monodromy cone associated to the degeneration $f\colon(X,L)\to\Delta^{k}$. This definition extends to any family of $g$-dimensional PPAVs $f^{\ast}\colon(X^{\ast},L^{\ast})\to(\Delta^{\ast})^{k}$ with unipotent monodromy.
@@
Note that $\widetilde{\Psi}$ descends because the $(\tau_{1}B_{1}+\cdots+\tau_{k}B_{k})$ term cancels the equivariance of $\widetilde{\Phi}$, and so makes the map invariant under the action of ${\mathbb{Z}}^{k}$. We now recall Schmid’s multivariable nilpotent orbit theorem [59, Thm. 4.12], applied to our setting:
@@
${\mathcal{P}}_{g}^{+}=\textstyle{\mathcal{P}}_{g}\sqcup\bigsqcup_{V_{1}}{\mathcal{P}}_{g-1}\sqcup\bigsqcup_{V_{2}}{\mathcal{P}}_{g-2}\sqcup\cdots\sqcup\bigsqcup_{V_{g}}{\mathcal{P}}_{0}$
where the $V_{i}\subset{\mathbb{Z}}^{g}$ range over all primitive integral sublattices of ${\mathbb{Z}}^{g}$ of dimension $i$, and the relevant copy of ${\mathcal{P}}_{g-i}$ is the cone of positive-definite bilinear forms on ${\mathbb{R}}^{g}/(V_{i}\otimes{\mathbb{R}})$. See Figure 22 to visualize the projectivization of ${\mathcal{P}}_{2}^{+}$, which is a cusped hyperbolic disk.
@@
We have that $W_{-2}^{\perp}=W_{-1}$ where the perpendicular is taken with respect to $L$. Thus $L$ descends to a unimodular symplectic form on $W_{-1}/W_{-2}$. In fact, by [59, Thm. 6.16], we have the following fundamental theorem on the existence of the limit MHS:
@@
with the $10$ singularities given by the $S_{6}$-orbit of the point $1:1:1:-1:-1:-1]\in{\mathbb{P}}^{5}$. Then, the universal deformation (whose existence follows from [30, Cor. 4.2], for example)
@@
see [34, Sec. 7.1.10], and generate a primitive integral Lagrangian subspace $W_{-2}\subset H_{3}(Y_{t},{\mathbb{Z}}(-1))$ of rank $5$. The Dehn twist about the three-sphere $\gamma_{i}$ gives the formula
@@ 2.4. Toric varieties
Then the gluing $Y({\mathfrak{F}})\coloneqq\bigcup_{\tau\in{\mathfrak{F}}}Y(\tau)$ of these affine schemes along the natural open immersions corresponding face inclusions gives the toric variety $Y({\mathfrak{F}})$, see [31, Sec. 1.4].
@@
As in the usual theory of toric varieties, the torus orbits of dimension $d$ in $Y({\mathfrak{F}})$, isomorphic necessarily to $({\mathbb{C}}^{*})^{d}$, correspond bijectively to cones of codimension $d$ in $\mathfrak{F}$}, see [31, Sec. 3.1].
@@
The toric variety $Y=Y_{H}$ associated to the polytope $H$ is the union of ${\operatorname{Spec}}\,{\mathbb{C}}[{\bf M}_{F}]$ ranging over all faces, see also [31, Sec. 1.5]. If, furthermore, all faces of $H$ are integral polytopes, then there is a canonically defined torus-equivariant holomorphic line bundle $\mathcal{L}$} on $Y$, given by gluing together line bundles on each affine chart ${\operatorname{Spec}}\,{\mathbb{C}}[{\bf M}_{F}]$, in a manner which locally agrees with the recipe in [31, Sec. 3].
@@ 2.5. Toroidal extensions
A fan $\mathfrak{F}$} for ${\mathcal{A}}_{g}$ is a rational polyhedral fan, whose support is contained in ${\mathcal{P}}_{g}^{+}\subset{\rm Sym}_{g\times g}({\mathbb{R}})$, and such that $\mathfrak{F}$} is $\operatorname{GL}_{g}({\mathbb{Z}})$-invariant under the action $B\mapsto ABA^{T}$, with finitely many orbits of cones.
@@
(8) $\overline{E}\colon U_{\mathbb{Z}}\backslash{\mathcal{H}}_{g}\hookrightarrow U_{\mathbb{Z}}\backslash{\rm Sym}_{g\times g}({\mathbb{C}})\simeq\operatorname{Sym}_{g\times g}({\mathbb{C}}^{\ast})\simeq{\rm Sym}_{g\times g}({\mathbb{Z}})\otimes{\mathbb{C}}^{*}\simeq U_{\mathbb{Z}}\otimes{\mathbb{C}}^{*}.$
@@
More generally, the same construction applies to any $\operatorname{GL}_{g}({\mathbb{Z}})$-invariant fan $\mathfrak{F}$}, and the resulting toroidal extension is relatively proper over the 0-dimensional cusp of the Baily-Borel [11] compactification $\overline{{\mathcal{A}}}_{g}^{\rm BB}={\mathcal{A}}_{g}\sqcup{\mathcal{A}}_{g-1}\sqcup\cdots\sqcup{\mathcal{A}}_{1}\sqcup{\mathcal{A}}_{0}$ if and only if ${\rm Supp}({\mathfrak{F}})={\mathcal{P}}_{g}^{+}$.
When $\mathfrak{F}$} contains cones supported in ${\mathcal{P}}_{g}^{+}\setminus{\mathcal{P}}_{g}$, the gluing defined by (10) further extends along the intermediate-dimensional strata of the Baily-Borel compactification. This is the toroidal extension ${\mathcal{A}}_{g}\hookrightarrow{\mathcal{A}}_{g}^{{\mathfrak{F}}}$ of the orbifold ${\mathcal{A}}_{g}$.
When ${\rm Supp}({\mathfrak{F}})={\mathcal{P}}_{g}^{+}$, we notate the toroidal extension by ${\mathcal{A}}_{g}\hookrightarrow\overline{{\mathcal{A}}}_{g}^{\mathfrak{F}}$; it is proper, and we call it a toroidal compactification. It follows from [9, Thm. 5.2] that ${\mathcal{A}}_{g}^{\mathfrak{F}}$ or $\overline{{\mathcal{A}}}_{g}^{\mathfrak{F}}$ is a DM algebraic stack, in the former case by refining and extending $\mathfrak{F}$} to a fan with full support ${\mathcal{P}}_{g}^{+}$.
@@
Next, it follows from the exponential convergence (2) of $\Phi$ towards $\Phi_{\rm nilp}$ (in the invariant metric on ${\mathcal{H}}_{g}$) that $\Phi\colon(\Delta^{*})^{k}\to{\mathcal{A}}_{g}$ admits a continuous extension $\Delta^{k}\to{\mathcal{A}}_{g}^{{\mathbb{B}}}$. Since $\Delta^{k}$ is normal, Riemann’s removable singularities theorem implies that this continuous extension is holomorphic. It is unique because the toroidal extension is a separated analytic stack. ∎
@@
Then, a toroidal extension of $\widetilde{{\mathcal{A}}}_{g}$ at this Baily–Borel cusp has the “advantage” that a fan need only be $\Gamma(n)$-invariant. Furthermore, the fundamental domain for the action of $\Gamma(n)$ is larger than the fundamental domain for that action of $\operatorname{GL}_{g}({\mathbb{Z}})$. In particular, given any polyhedral cone ${\mathbb{B}}\subset{\mathcal{P}}_{g}^{+}$, there exists some $n$ (depending on $\mathbb{B}$}) for which $\mathbb{B}$} embeds into the quotient $\Gamma(n)\backslash{\mathcal{P}}_{g}^{+}$. The preceding results also apply at this cusp, since $\Gamma(n)\cdot{\mathbb{B}}$ now defines a fan.
@@
To verify that the monodromy representation of $f^{*}$ is as specified, consider a $1$-parameter family $\Delta^{*}\to Y^{*}$ which extends to a map $\Delta\to Y$ sending $0\mapsto 0$. The monodromy over $\Delta^{*}$ can be computed in the étale-local model as the monodromy of the family of PPAVs over a co-character. Such co-characters correspond to lattice points, in $\tau\cap{\mathbb{Z}}^{k}$. Since $f^{*}\colon X^{*}\to Y^{*}$ is pulled back from $V^{*}$, the monodromy representation is pulled back along the morphism of cocharacter lattices, given by the inclusion $\Lambda\hookrightarrow{\rm Sym}_{g\times g}({\mathbb{Z}})$ which, in particular, restricts $\phi\colon{\mathbb{Z}}^{k}\to{\rm Sym}_{g\times g}({\mathbb{Z}})$. Finally, it suffices to observe that the monodromy over the cocharacter $B\otimes{\mathbb{C}}^{*}\subset{\rm Sym}_{g\times g}({\mathbb{Z}})\otimes{\mathbb{C}}^{*}$ is canonically identified with $B\in{\mathcal{P}}_{g}^{+}\cap{\rm Sym}_{g\times g}({\mathbb{Z}})$, see e.g. (12).